40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Thu Dec 9 23:17:47 2010, kiwamu, Update, SUS, beam pointing has been done 
    Reply  Fri Dec 10 09:58:57 2010, Aidan, Update, SUS, beam pointing has been done 
       Reply  Sat Dec 11 00:41:55 2010, kiwamu, Update, IOO, alignment of IP_ANG is done 
          Reply  Tue Dec 14 01:04:23 2010, kiwamu, Update, SUS, alignment of ITMs and PRM done 
             Reply  Tue Dec 14 15:15:19 2010, kiwamu, Update, SUS, put ETMY and aligned the beam 
                Reply  Tue Dec 14 22:42:56 2010, kiwamu, Update, SUS, installed the last TT . And some more progresses 
                   Reply  Thu Dec 16 01:23:43 2010, kiwamu, Update, SUS, SRM OSEMs installed DSC_2757_ss.jpg
Message ID: 4063     Entry time: Thu Dec 16 01:23:43 2010     In reply to: 4055
Author: kiwamu 
Type: Update 
Category: SUS 
Subject: SRM OSEMs installed 

  [Zach and Kiwamu]

 We worked on some more vacuum businesses.  Today we finished did the following works:

 - alignment of the POX mirrors 

 - alignment of the POP1 and POP2 mirrors

 - installation of OSEMs onto SRM

 - alignment of the SRM tower

 

 (alignment of POP mirrors)

 Since a beam on the POP path was quite too weak to see even by IR viewers, we used a He-Ne laser to imitate the real beam instead.

We injected the He-Ne beam from an optical bench to the chamber, and made it go through the PRM and PR2 by using some steering mirrors.

 

(OSEM cables)

 The pin assignment was flipped in a way of mirror image due to the extension cables which cause a mirroring.

 So we made mirroring connectors to flipp them back to the correct pin assignment, and plugged the mirroring connectors in between the feedthrough of the BS chamber and the SRM satellite box.

This is a picture showing how they are connected now.

DSC_2757_ss.jpg

ELOG V3.1.3-